A metallic wire of density d floats horizontal in water. The maximum radius of the wire so that the wire may not sink will be?  ( surface tension of water is T and angle of contact is 0)

Grade:12

1 Answers

Arun
5 years ago
Dear student 
We know the SURFACE TENSION as:
T = force/length 
or  = F/l
F =>  mg = g*PI*r^2*d*length 
calculating r value from here we get, 
r = root( m / g*pi*d*length)
 
 
Regards
Arun